OpenStudy (nicky):

What is $699.99 in percent

OpenStudy (nicky):

@ganeshie8

OpenStudy (nicky):

@ParthKohli

OpenStudy (nicky):

@sammixboo

OpenStudy (nicky):

@sweetburger

OpenStudy (nicky):

MEDAL+FAN

OpenStudy (nicky):

I really need help

OpenStudy (blurbendy):

$699.99 is simply a dollar amount. To get a percentage, you'd need another number. Are you sure you're not missing something?

OpenStudy (nicky):

I know you can't put money into a percent, but how can you put it into a decimal, then a percent? if i understand this is what we need to do

OpenStudy (nicky):

oh, well I also have $5.49

OpenStudy (blurbendy):

Well, I guess you could say 25 cents = .25 = 25/100 = 1/4 one dollar would be 1/1 = 1.00 = 100/100 = 1 So, what would $5.49 be?

OpenStudy (nicky):

wait a sec.

OpenStudy (nicky):

699.99? I am stupid idk is that right?

OpenStudy (blurbendy):

No, 5.49 would just be 549/100 = 549 percent

OpenStudy (blurbendy):

so, what would 699.99 convert to?

OpenStudy (nicky):

69999/100=

OpenStudy (blurbendy):

yes, which would be what percent?

OpenStudy (nicky):

wait a sec. my computed is bogging

OpenStudy (nicky):

idk

OpenStudy (mrnood):

it is normal to express a number as a percentage of another but basically 'percent' means 'divide by 100' so 699 = 69900/100 = 69900 percent do the same with YOUR number...

OpenStudy (nicky):

sorry my computer is having a hard time....

OpenStudy (nicky):

it close my tables ahhh

OpenStudy (blurbendy):

If you don't understand bigger numbers, start small first. Do you what $2.00 is in percent form based on what you know?